C=500+0.4yd,MPs=0.6,I=500,g=700,x=200 and m=500+0.2y the tax is 20%of y the population is 90 people calculate in equilibrium


1
Expert's answer
2021-09-22T07:50:49-0400

Solution:

At equilibrium: Y = AE

Y = C + I + G + X – M

T = 20"\%" of Y = 0.2Y

C = 500 + 0.4Yd = 500 + 0.4(Y – T) = 500 + 0.4(Y – 0.2Y)

I = 500

G = 700

X = 200

M = 500 + 0.2Y

Y = 500 + 0.4(Y – 0.2Y) + 500 + 700 + (200 – 500 + 0.2Y)

Y = 500 + 0.32Y + 500 + 700 + (200 – 500 + 0.2Y)

Y = 500 + 500 + 700 + 200 – 500 + 0.32Y – 0.2Y

Y = 1400 + 0.12Y

Y – 0.12Y = 1,400

0.88Y = 1,400


Y = 1,591


The equilibrium income (Y) = 1,591